1. consider the following information for a hypothetical economy:


C = 200 + 0.75(Y – T)        I=G=250    T=200    where, Y=C+I+G

All figures are in millions except the measures of responsiveness.

  1. What is the value of marginal propensity to save(MPS) ?Interpret it.
  2. Find the equilibrium level of national income
  3. Find the government expenditure multiplier ? using this multiplier , find the effect of 50 million increase in government purchase on equilibrium income ?
1
Expert's answer
2022-03-14T13:15:09-0400

Solution

"C = 200 + 0.75(Y \u2013 T)"

"I=G=250"

"T=200"

Therefore;

 "Y=C+I+G"

"Y=200+0.75(Y-200)+250+250"

"Y=200+0.75Y-150+500"

"Y-0.75Y=550"

"0.25Y=550"

"Y=\frac{550}{0.25}"

"Y=2,200"


a)Marginal propensity to save="\frac{change in savings}{change in disposable income}"

"MPS=1-0.75"

Mps is 0.25 of the total income

Therefore;

"MPS=0.25"


b) Equilibrium

"Y=2,200"


c)Government expenditure multiplier

"GEM=\frac{change in income}{change in government spending}"


"GEM=\frac{2,200}{250}"


"=8.8"


Increase of 50 million will be


"=\frac{8.8\times50,000,000}{100} +50,000,000"


"=54,400,00"